# Short Film Dataset (SFD): A Benchmark for Story-Level Video Understanding

### v1

@@ 5 Discussion
We introduced the Short Film Dataset (SFD), a long-form video understanding benchmark featuring complete and narrative-driven amateur short movies. It includes two types of QA tasks, multiple-choice and open-ended, that are automatically generated by LLMs and manually curated by human annotators. Compared to other long-form video datasets, SFD stands out by offering richer, story-oriented content with tasks specifically tailored for long-term reasoning, as validated by our experiments. Unlike most movie datasets that use copyrighted commercial films prone to data leakage with LLMs, SFD relies on amateur films, which are publicly accessible and have a limited online presence. Furthermore, our analysis indicates that while language-based methods achieve performance levels comparable to humans, state-of-the-art vision-based and multimodal methods fall behind human evaluation and they still have considerable room for improvement, highlighting the need for further advancements. In conclusion, we believe that our SFD paves the way for accessible, and comprehensive long-term movie understanding that is not known a priori by LLMs, thus helping the community develop robust methods. Our dataset and code will be made publicly available.
6 Acknowledgement
This work was supported by ANR-22-CE23-0007, Hi!Paris grant and fellowship, and was granted access to the High-Performance Computing (HPC) resources of IDRIS under the allocations 2023-AD011014489 made by GENCI. We would like to thank Nicolas Dufour, Robin Courant, Pierre Vassal, and the anonymous reviewers for their insightful comments and suggestions. We also express our gratitude to the annotators of the human study for their help.
@@ 6 Acknowledgement
In this appendix, we show additional materials of the paper Short Film Dataset (SFD): A Benchmark for Story-Level Video Understanding. We present the following items:
Appendix A Checklist
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
Did you describe the limitations of your work? [Yes]
Did you discuss any potential negative societal impacts of your work? [Yes] See Supplementary Materials Section B
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] See Supplementary Materials Section B
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A]
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] https://huggingface.co/datasets/rghermi/sfd
- (b)
Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] Our experiments consist of benchmarking various aspects of our dataset. For a given compute, instead of performing one experiment several times (with different random seeds), we chose to experiment with several different multimodal methods and multiple LLMs. We observed that they all report similar performances and overall follow the same trend; hence, we argue that our findings are generalizable.
Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 3.3
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
If your work uses existing assets, did you cite the creators? [Yes] See Section B.
Did you mention the license of the assets? [Yes] See Section B.
Did you include any new assets either in the supplemental material or as a URL? [N/A]
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Section B.
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Section F.
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix B Ethical considerations
@@
We do not distribute raw video content; instead, we provide only URLs redirecting to YouTube, where the full copyright and licensing rights of creators are acknowledged. Additionally, by sharing the URLs, YouTube’s Terms of Service regarding exclusive rights are respected.333https://www.youtube.com/static?template=terms (Section: License to Other Users).
The dataset will be used solely for academic research purposes. To further protect the metadata and other information, we apply the CC BY-NC-SA 4.0 (Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International) license to our dataset.
@@
Impact on creative process: Progress on movie understanding may encourage filmmakers to tailor their work to optimize for what is favoured by algorithms, potentially stifling creative freedom and depriving innovation.
Appendix C Data Preprocessing and Statistics
SFD | |
#Movies | 1,078 |
#Clips | 161,321 |
#Clips / Movie | 151 |
Avg. Clip Len. (s) | 5.43 |
#Captions | 161,321 |
#Location Tags | 161,321 |
SFD | |
#Face Instances | 56,265 |
#Unique Faces | 2,598 |
#Unique Faces / Movie | 2.41 |
#Subtitles | 23,820 |
SFD, alongside the videos and subtitles (see Section 2), also offers a comprehensive set of annotations: clips (shots), shot captions and scales, location tags, and face tracks. These resources, together with tools for downloading, preprocessing, and annotation, will be made available as part of the dataset release. Below, we detail their creation processes.
Clips (Shots). In the context of movies, a clip corresponds to a shot, characterized by its composition, length, and camera movement. We use PySceneDetect444https://www.scenedetect.com/ to segment movies in SFD into 161,321 shots, with each movie containing an average of 151 shots, as shown in Table 5(a). The number of shots varies by genre; for example, comedies average 163 shots, while dramas have about 146.
Clip information. For each shot, we extract detailed shot-level information. We generate shot captions using image and video captioning tools (LLaVA [36], BLIP-2 [30], Video-LLaVA [33]), and we identify locations using CLIP features [43] based on a curated list of locations from MovieCLIP [6].
Face Tracks. We use an advanced facial recognition pipeline 555https://github.com/serengil/deepface to identify over 56k+ face instances in the dataset, corresponding to 2.6k+ unique characters, with an average of 2.41 characters per movie, as shown in Table 5(b).
These annotations enrich the dataset, providing deeper insights for tasks related to character analysis and interaction within the films.
@@ Appendix C Data Preprocessing and Statistics
In Figure 8, we perform a word cloud analysis for three video domains commonly used in the literature: instructional videos [71], egocentric videos [39], and movies (ours). We notice that instructional and egocentric videos are focused on actions (verbs like make, go, cut, use, pick, put) and objects (e.g. table, sugar, oil, thing, stuff, table, paper), with an emphasis on body parts for egocentric videos (right hand, left hand). For movies, the vocabulary is more focused on characters (e.g. man, woman, boy, girl, young, old), relationships (e.g. mother, father, family, couple), time (e.g. day, night, year), and locations (e.g. home). These clearly show the story component of SFD.
Appendix D Technical details on baselines
In this section, we present additional details and results for the baselines discussed in Section 3.3 and Table 3 main paper, where we report the benchmark performance of the state-of-the-art video understanding methods on our SFD.
FrozenBiLM [5]. We use FrozenBiLM pre-trained on WebVid10M[5]. Following[39], we reformulate the question-answering task for this model with the following prompt: ‘Question: {question} Is it {option_i}?’. We get logits for the yes/no tokens for each option and take the maximum as the final prediction. We report results with and without subtitles. We compute the accuracy score, by comparing the predicted option index and the correct option index. As it is a masked language model, FrozenBiLM is not suited to open-ended question answering.
mPLUG-Owl2 [74]. We use the instruction-tuned version (‘MAGAer13/mplug-owl2-llama2-7b’), based on ViT-L and LLaMa-7b. Inference is the same as FrozenBiLM. In the open-ended question-answering setting, we only provide the question and record a free-form text output: ‘Question: {question} Please answer with a short sentence.’. Following [37], we use GPT-3.5 (‘gpt-3.5-turbo’) to compute the similarity between the predicted and the correct answer. We report this LLM-based accuracy.
Video-LLaVA [33]. We use the fine-tuned version (‘LanguageBind/Video-LLaVA-7B-hf’), based on LanguageBind and Vicuna-13B v1.5. Inference is the same as mPLUG-Owl2, in both multiple-choice and open-ended question-answering settings.
LLoVi [79]. Following the official implementation, we use LLaVA (‘llava-hf/llava-1.5-13b-hf’) to extract captions at 0.5 FPS and GPT-3.5 (‘gpt-3.5-turbo’) as a backbone LLM. We experiment with the prompting strategy proposed in the original paper [79]. The prompt for multiple-choice question answering is as follows:
‘Please provide a single-letter answer (A, B, C, D, E) to the following multiple-choice question, and your answer must be one of the letters (A, B, C, D, or E). You must not provide any other response or explanation. You are given some language descriptions of a video. The video is {duration} seconds long. Each sentence describes a {clip_length}s clip. The descriptions are sequential and non-overlapping which cover the whole video exactly. Here are the descriptions: {captions}. You are going to answer a multiple-choice question based on the descriptions, and your answer should be a single letter chosen from the choices. Here is the question: {question}. Here are the choices. A: {option_0} B: {option_1} C: {option_2} D: {option_3} E: {option_4}")’
LangRepo [26]. Similar to LLoVi, LangRepo is another LLM-backed video understanding method. We use the open-sourced official implementation, Mixtral 87B (‘mistralai/Mixtral-8x7B-Instruct-v0.1’). Unlike LLoVi, LangRepo operates in two stages: The first is summarization, where LangRepo consolidates similar information from descriptions, captions, subtitles, or both, using specific prompts following the official implementation:
‘You are a helpful expert in video analysis. You are given a list of {num_to_rephrase} language descriptions for a video. Each sentence describes a {clip_length}s clip. Here are the descriptions as a list:{memory}. Please summarize and rephrase each item in the list as a single sentence of {num_words_in_rephrase} words. Keep the same original subject. Keep all information intact without leaving anything out. Return only the rephrased list of {num_to_rephrase} descriptions in the same order, without additional details.’
In the second stage, the question-answering is based on the rephrased summarization with prompts:
‘Here are the descriptions: {narration}. You are going to answer a multiple-choice question based on the descriptions. Here is the question: {question}? You must select one of these choices as the answer: A: {option_0} B: {option_1} C: {option_2} D: {option_3} E: {option_4} The correct answer is:’
MovieChat [54]. We use the official implementation of MovieChat (‘rese1f/MovieChat’, which incorporates LLaMa2-7B (‘meta-llama/Llama-2-7b’ and corresponding Video-LLaMa-7B (‘DAMO-NLP-SG/Video-LLaMA-2-7B-Finetuned’). Similarly, we use GPT-3.5 (‘gpt-3.5-turbo’) [42] to compute the accuracy of responses. Following the official implementation, we use the following prompt:
‘By watching the video and combining it with the caption/subtitle from video , you are required to answer a question with multiple options: Question: {question}; Option A: {option_0}; Option B: {option_1}; Option C: {option_2}; Option D: {option_3}; Option E: {option_4}; Answer to the multiple-choice question based on the textual information extracted from the movie. If you don’t know the answer, please output ‘Unknown’. Your output should be just one of A, B, C, D, E and nothing else.’
TimeChat [47]. We use the official implementation of TimeChat (‘ShuhuaiRen/TimeChat-7b’). Following the official implementation, the prompt and the accuracy computation details are the same as those in the previous MovieChat model.
For all models, we also incorporate subtitles as input. For that, we concatenate each query with the following text: ‘Subtitles: {subtitles}’. Subtitles are formatted as follows: ‘{start_timestamp} - {end_timestamp} {subtitle_i}’. We trim subtitles if they cannot fit into the input size. Results and analysis can be seen in Section 3.3 and Table 3.
Appendix E Technical details on the temporal window study
In this section, we present additional details and results for the temporal window study (see Section 3.2 and Figure 7). The goal of the temporal window study is to evaluate the performance of various methods on SFD at different temporal levels—shot, scene, and movie—across all modalities. Along with the discussed FrozenBiLM and LLoVi, we examine Video-LLaVA and LangRepo.
At the shot-level, we truncate the input prompt to include only information from a specific shot. For FrozenBiLM and Video-LLaVA, it means that we only include subtitles and frames extracted from the shot. For LLoVi and LangRepo, we only include subtitles and captions extracted from the shot. Hence, we get logits for each answer option and shot, aggregated in a matrix of shape : we obtain a final prediction by taking the maximum logit over both dimensions. Note that the shot-level inference takes too much time to compute for LangRepo ( A100 GPU-hours per experiment), as this method involves many steps of rephrasing and summarizing.
The scene-level inference is similar to the shot-level one, the main difference being that we constrain the input information (subtitles and frames/captions) to the scene-level, which arbitrarily corresponds to 10 shots in our case. We end up with a logit matrix of size .
The movie-level inference is straightforward: it is the same as in Section D.
Vision | Language | Vision-Language | ||
FrozenBiLM | Shot-Level | 22.1 | 30.9 | 30.8 |
Scene-Level | 22.7 | 34.8 | 34.8 | |
Movie-Level | 23.4 | 38.2 | 38.6 | |
Video-LLaVA | Shot-Level | 16.0 | 24.4 | 24.0 |
Scene-Level | 23.7 | 20.1 | 20.1 | |
Movie-Level | 34.2 | 21.3 | 24.7 | |
LLoVi | Shot-Level | 32.8 | 38.5 | 50.1 |
Scene-Level | 34.2 | 51.2 | 55.4 | |
Movie-Level | 30.8 | 64.2 | 55.6 | |
LangRepo | Shot-Level | - | - | - |
Scene-Level | 26.9 | 27.9 | 28.1 | |
Movie-Level | 29.0 | 32.1 | 31.0 | |
Human | Movie-Level | 59.0 | 70.9 | 89.8 |
The numerical results of the temporal window study are reported in Table 6.
Along the temporal direction (i.e. row-wise for each method), we observe that: (i) generally, all methods exhibit consistent gains with larger temporal windows confirming the importance of long-term understanding for our designed questions, and (ii) exceptions are noted, such as with Video-LLaVa in Language and Vision-Language settings, likely due to its low precision (near random chance at 20%) which diminishes the impact of the temporal window.
For different modalities (i.e. column-wise), we make the following observations: (i) Corroborating the argument in the main paper (Section 3.3), the Language-only models, in general, achieve better performances, e.g. LLoVi reaches the best performance at movie-level in the language-only setting (64.2%). (ii) The vision-only setting results in the weakest performance, revealing the largest gap compared to human performance (last row). This suggests significant potential for improvement in visual models. (iii) Current modality-fusion methods show limited effectiveness. VLM models like FrozenBiLM and Video-LLaVA only demonstrate marginal improvements when fusing vision and language, with performance increment from 38.2% to 38.6% for FrozenBiLM and from 21.3% to 24.7% for Video-LLaVa. However, all LLM-based methods experience a regression after modality fusion.
Overall, the human benchmark highlights that there remains a substantial gap between the best methods and human performance, particularly in vision-language integration over extended temporal windows.
Appendix F Human Study
@@
In this section, we present additional details for the conducted human study (see Section 3.3, ‘User Study’). We conducted a comprehensive human study to evaluate the upper-limit performance of our benchmark. Participants were asked to watch a movie and subsequently answer multiple-choice questions based on the content. To analyze the impact of different modalities, we implemented two ablation conditions: a muted version, where participants had access to the vision-only information without audio or subtitles, and a blind version, where participants were provided with the audio-only information, in the form of textual subtitles, without any video.
@@
To avoid any bias, we ensured that different sets of participants were assigned to the full, vision-only, and audio-only settings, preventing any overlap and potential influence from experiencing multiple conditions. Movies were randomly assigned to users to ensure a diverse range of data. The entire study was facilitated through an integrated interface, which restricted access to external information such as movie loglines and synopses, with only the movie title being displayed (see Figure 9).
Appendix G Technical details on data leakage experiment
Name | SFD | MovieQA | LVU |
@@ Appendix G Technical details on data leakage experiment
In this section, we include additional information on the data leakage study mentioned in the main paper Section 3.1. The objective of our data leakage study is to demonstrate that current large language models (LLMs) possess prior knowledge about commercial movies. This is because LLMs have been exposed to vast amounts of movie-related textual information during pretraining, including synopses, blog posts, news articles, and more. The experiment is set up as an open-ended question-answering task where the input consists of a movie title and a corresponding question, and the output is a free-form text generated by the LLM. We then evaluate the answer using an LLM to compute the similarity between the predicted and the correct answer, following [37].
We choose ten of the most recent LLMs, among which six are open-source models: Gemma 2B, Gemma 7B, LLaMA-3 8B, LLaMA-3 70B, Mistral 7B, Mixtral 8x7B; and four commercial models: Claude 3 Haiku, Claude 3 Sonnet, GPT-3.5, GPT-4. The open-source models are run using the Hugging Face transformers library 666https://github.com/huggingface/transformers. The commercial models are run using their official API: the OpenAI API 777https://openai.com/api for GPT and the Anthropic API 888https://www.anthropic.com/api for Claude 3.
We choose three movie datasets: MovieQA [56], LVU [64], and SFD (ours). For MovieQA and SFD, the question-answer pairs are directly available. In the case of LVU, we re-formulated the classification tasks into templated questions to fit the open-ended question-answering format. For example, the director classification task is converted into the question ‘Who directed the movie {movie_title}?’.
The detailed results of the data leakage experiment are presented in Table 7 and Figure 10. Their analysis shows that LLMs exhibit significantly higher data leakage on MovieQA and LVU datasets compared to SFD. Models like GPT-4 and Claude 3 Sonnet answer correctly to more than 70% of the questions of both datasets, without any context except the movie title. The scatter plot emphasizes a correlation between data leakage and model knowledge (i.e. MMLU performance), with higher MMLU scores correlating with increased leakage issues on compromised datasets. This trend is consistent across both open-source and commercial models, highlighting the vulnerability of larger and more advanced models to memorizing training data.
In sharp contrast, the relatively low and flat performance trend on SFD suggests it is a more robust benchmark for evaluating model capabilities. These findings emphasize the need for robust evaluation datasets to accurately assess model performance and mitigate the impact of data leakage.
Appendix H Samples from SFD
In the following pages, we show some samples from SFD. For each sample, we display the title, the logline, the synopsis, a few frames, some subtitles, and the designed multiple-choice questions. See Figures 11, 12,13, 14.
Appendix I Prompts

### v2

@@ 5 Discussion
We introduced Short-Films 20K (SF20K), a long-form video understanding dataset featuring 20,143 amateur movies. It includes two QA tasks, multiple-choice and open-ended, automatically generated by LLMs and, in the test set, manually curated by human annotators. Compared to other video datasets, SF20K stands out by offering richer, story-oriented content with tasks specifically tailored for long-term reasoning, as validated by our experiments. Unlike other movie datasets that use copyrighted commercial movies prone to data leakage with LLMs, SF20K relies on amateur films, which are publicly accessible and have limited online presence. Our analysis indicates that state-of-the-art multimodal methods fall behind human accuracy, with performances being mostly driven by the language modality (i.e. subtitles), implying rooms for improvement and the need for further advancement of vision-language models. Finally, we show that instruction-tuning VLMs on SF20K-Train improves performance across modalities and tasks. In conclusion, we believe that SF20K paves the way for accessible and comprehensive movie understanding that is not known a priori by LLMs, thus helping the community to develop robust methods.
6 Acknowledgments
This work was supported by ANR-22-CE23-0007, Hi!Paris grant and fellowship, and was granted access to the High-Performance Computing (HPC) resources of IDRIS under the allocations 2023-AD011014489 made by GENCI. We would like to thank Nicolas Dufour, Robin Courant, and Pierre Vassal for their insightful comments and suggestions. We also express our gratitude to the annotators of the human study for their help.
@@ 6 Acknowledgments
In the following, we present the following items:
Ethical considerations
Data pre-processing and statistics
Technical details on the data leakage experiment
Human study
Technical details on baselines
Failure cases analysis
Technical details on the temporal window study
Prompts
Samples
Datasheet for datasets
Appendix A Ethical considerations
@@
We do not distribute raw video content; instead, we provide only URLs redirecting to YouTube and Vimeo, where the full copyright and licensing rights of creators are acknowledged. Additionally, by sharing the URLs, YouTube’s and Vimeo’s Terms of Service regarding exclusive rights are respected.222https://www.youtube.com/static?template=terms (Section: License to Other Users), https://vimeo.com/terms (Section: Licenses Granted by You).
The dataset is intended exclusively for academic research purposes. To ensure the protection of the dataset and all its associated elements, we use the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International (CC BY-NC-SA 4.0) license.
@@
Impact on creative process: Progress on movie understanding may encourage filmmakers to tailor their work to optimize for what is favoured by algorithms, potentially limiting creative freedom and depriving innovation.
Appendix B Data pre-processing and statistics
SF20K provides not only videos but also a comprehensive set of annotations to facilitate deeper analysis and understanding. These annotations include subtitles, shot boundaries, keyframe captions, and face tracks. Together with tools for downloading, pre-processing, and annotating the data, these resources are released as part of the dataset. Below, we detail the processes involved in creating these annotations.
| #Movies | #Hours | #QA Pairs | Source | Annotation | |
| SF20K | 20,143 | 3,584 | 191,007 | Subtitles | Auto |
| SF20K-Test | 1,072 | 244 | 4,885 | Synopsis | Auto + Manual |
| SF20K-Test-Silent | 90 | 20 | 419 | Synopsis | Auto + Manual |
| SF20K-Test-Expert | 50 | 11 | 570 | Movie | Manual |
| SF20K | |
| #Movies | 20,143 |
| #Shots | 2,396,583 |
| #Shots / Movie | 119 |
| Avg. Shot Len. (s) | 5.52 |
| #Shot Captions | 2,396,583 |
| #Subtitles | 1,088,140 |
| #Face Tracks | 1,083,850 |
Shot boundaries. In the context of movies, a shot is a continuous sequence of frames captured without interruption by a single camera, often representing a single perspective or action within the scene. This corresponds to a clip in other video datasets. We use PySceneDetect333https://www.scenedetect.com/ with a threshold equal to 3.0 to segment the 20,143 movies in SF20K into 2,396,583 shots, with each movie containing an average of 119 shots.
Keyframe captions. For each shot, a detailed caption is generated by extracting the middle frame of the shot and processing it with Llama-3.2-Vision [43]. The model is specifically prompted to identify and describe relevant details, including the time, location, character appearances, mood, and overall atmosphere. We end up with one caption per shot, i.e. 2,396,583 captions, for a total of 59,760,112 words. These textual annotations provide a good basis for further work on video generation models.
Subtitles. We extract missing subtitles from the audio with Whisper Large-v3 Turbo [47], using the translation mode. We end up with 1,088,140 dialog sentences, for a total of 12,038,891 words.
Face tracks. We use a facial recognition pipeline 444https://github.com/serengil/deepface to identify over 1,083,850 face tracks. We first detect faces with RetinaFace [14], and then compute face tracks by matching face detections among consecutive frames of the same shot with IoU matching. These annotations enrich the dataset, providing deeper insights for tasks related to character analysis and interaction within films.
@@ Appendix B Data pre-processing and statistics
Word cloud analysis. In Figure 8, we perform a word cloud analysis for three video domains commonly used in the literature: instructional videos [77], egocentric videos [41], and movies (ours). We notice that instructional and egocentric videos are focused on actions (verbs like make, go, cut, use, pick, put) and objects (e.g. table, sugar, oil, thing, stuff, table, paper), with an emphasis on body parts for egocentric videos (right hand, left hand). For movies, the vocabulary is more focused on characters (e.g. man, woman, boy, girl, young, old), relationships (e.g. mother, father, family, couple), time (e.g. day, night, year), and locations (e.g. home). These differences highlight the story-centric nature of the SF20K dataset.
Subset statistics. In Table 7, we provide the key statistics of SF20K and its subsets. One key element to highlight is the annotation process. SF20K comes with almost 200K question-answer pairs, used for instruction tuning. Those have been generated automatically from subtitles. On the contrary, SF20K-Test questions have been automatically generated from human-written synopses, and subsequently manually curated. This applies also to SF20K-Test-Silent, which is simply a subset of SF20K-Test with only silent movies. Finally, SF20K-Test-Expert is a fully manual test set, where questions are designed by annotators who watched the movie.
Question types. We use an LLM to automatically classify the question into 4 categories: setting-, character-, story-, and theme-related questions. In Table 10, we provide a few examples of questions for each question type from SF20K-Test.
Appendix C Technical details on the data leakage experiment
| Template: "In [MOVIE TITLE], [QUESTION]" |
| In ’Titanic’, what happened to the Heart of the Ocean? |
| In ’The Godfather’, how does Vito die? |
| In ’Harry Potter and the Chamber of Secrets’, who is Dobby’s master? |
| Question | Answer |
| Setting | |
| Where do John and Emma have their conversation? | On the bus. |
| Which city is the backdrop? | Liverpool. |
| What decade does the movie take place in? | 1990s. |
| Character | |
| Who is the main character? | Josh. |
| What is the nationality of the soldier who becomes a prisoner? | German. |
| How is Mr. Jones described? | As stern and harsh. |
| Story | |
| Can you name a specific tactic Tommy uses against Tiny? | He steals Tiny’s food at dinner and distracts her at a school-wide relay race. |
| What is the imminent threat to Greg’s residence? | The house is in foreclosure and the bank is about to repossess it. |
| What triggers Carla to believe her husband Frank will surprise her for their anniversary? | A bouquet left on her doorstep. |
| Theme | |
| What is a major theme? | The theme revolves around reconnecting and the complexities of honesty between exes. |
| What main theme is explored through Sarah’s experience? | The theme of romantic expectations versus present realities. |
In this section, we include additional information on the data leakage study mentioned in Section 4.1. The objective of our data leakage study is to demonstrate that current large language models (LLMs) possess prior knowledge about commercial movies. This is because LLMs have been exposed to vast amounts of movie-related textual information during pretraining, including synopses, blog posts, news articles, and more. The experiment is set up as an open-ended question-answering task where the input consists of the movie title and corresponding question (see Table 9 for a few examples), and the output is a free-form text generated by the LLM. We evaluate the answer using an LLM to compute the similarity between the predicted and the correct answer, following [39].
We choose ten of the most recent LLMs, among which six are open-source models: Gemma-2B, Gemma-7B, LLaMA-3-8B, LLaMA-3-70B, Mistral-7B, Mixtral-8x7B; and four commercial models: Claude 3 Haiku, Claude 3 Sonnet, GPT-3.5, GPT-4. The open-source models are run using the Hugging Face transformers library 555https://github.com/huggingface/transformers. The commercial models are run using their official API: the OpenAI API 666https://openai.com/api for GPT and the Anthropic API 777https://www.anthropic.com/api for Claude 3.
We choose three movie datasets: MovieQA [61], LVU [70], and SF20K-Test (Ours). For MovieQA and SF20K-Test, the question-answer pairs are directly available. In the case of LVU, we re-formulate the classification tasks into templated questions to fit the open-ended question-answering format. For example, the director classification task is converted into the question ‘Who directed the movie [MOVIE TITLE]?’.
The detailed results of the data leakage experiment are presented in Table 11. We notice that LLMs exhibit significantly higher data leakage on MovieQA and LVU datasets compared to SF20K. Models like GPT-4 and Claude 3 Sonnet answer correctly to more than 70% of the questions of both datasets, without any context except the movie title, and in an open-ended question-answering setup.
Appendix D Human Study
@@
In this section, we present additional details for the conducted user study (see Section 4.2). We conducted a comprehensive user study to evaluate the upper-bound performance on our benchmark. Participants were asked to watch a movie and subsequently answer multiple-choice questions based on the content. To analyze the impact of different modalities, we implemented two ablation conditions: a muted version, where participants had access to the vision-only information without audio or subtitles, and a blind version, where participants were provided with the audio-only information, in the form of textual subtitles, without any video.
@@
To avoid bias, we ensured that different sets of participants were assigned to different settings (full, vision-only, and audio-only), preventing any overlap and potential influence from experiencing multiple conditions. Movies were randomly assigned to users to ensure a diverse range of data. The entire study was facilitated through an integrated interface, which restricted access to external information such as movie loglines and synopses, with only the movie title being displayed (see Figure 9).
| Name | SF20K-Test | MovieQA | LVU |
@@
| Video ID | Question | Answer | Prediction |
| Visual cues | |||
| IehPbS9fbPY | How does Ernie’s suicide attempt fail? | The attempt is unsuccessful and results in a hole in his ceiling. | Ernie’s suicide attempt fails when he is interrupted by a significant moment of connection or realization, leading him to reconsider his actions. |
| xErsPPWDDdk | How do ordinary objects appear to Yumi during her visions? | Ordinary objects appear as ladles to Yumi during her visions. | I’m unable to provide specific details … |
| Character identification | |||
| wD–FpKwnvA | Who does Corey live with? | She lives with her grown daughter. | Corey lives with his mother, Karen. |
| mhxEnIp7YIg | Who invites Jane to the boxing gym? | Tisha | I’m unable to identify … |
| Temporal understanding | |||
| 0PKtGVJEpdg | Where does Becca encounter the man again after the incident? | She meets him at a bar one evening. | Becca encounters the man again after the incident in an alley. |
| 3PejXWtS4xs | What is Harris doing in the apartment at the beginning of the movie? | Waiting for his target to arrive. | At the beginning of the movie, Harris is in the apartment, having a conversation on the phone. |
Appendix E Technical details on baselines
In this section, we present additional details for the baselines discussed in Section 4.3, where we report the benchmark performance of the state-of-the-art video understanding methods on SF20K-Test. We detail the models evaluated, the prompting strategy, and the metrics used.
Models. We evaluate eleven open-source models: FrozenBiLM [6], mPLUG-Owl2 [80], TimeChat [52], Video-Llava [36], MovieChat [59], Llava-OneVision [32], Long-Llava [67], LongVA [84], Llava-Next-Video [85], Llama-3.2-Vision [43], and Pixtral [3]; and two commercial models: LloVi [83] (based on GPT-3.5 [45]), and GPT-4o-mini [46]. Most open-source models are available on Hugging Face’s transformers library 888https://huggingface.co/docs/transformers, except for FrozenBiLM 999https://github.com/antoyang/FrozenBiLM, mPLUG-Owl2 101010https://github.com/X-PLUG/mPLUG-Owl, and LongVA-7B 111111https://github.com/EvolvingLMMs-Lab/LongVA. Commercial models are available through their official API. Note that commercial models are not free; therefore, we are limited in their usage. We are only able to evaluate GPT-4o-mini on 10% of SF20K-Test (i.e. approximately 500 questions).
Prompting. We prompt all the models with the same queries, inspired by Gemini [51]. For multiple-choice question answering, we use the following prompt:
‘You will be given a question about a movie. Try to answer it based on the subtitles and the frames from the movie. Question: [QUESTION] Possible answer choices: (1) [OPTION 1] (2) [OPTION 2] (3) [OPTION 3] (4) [OPTION 4] (5) [OPTION 5] Output the final answer in the format "(X)" where X is the correct digit choice. DO NOT OUTPUT with the full answer text or any other words.‘
For open-ended question answering, we use the following prompt:
‘You will be given a question about a movie. Try to answer it based on the subtitles and the frames from the movie. Question: [QUESTION] Answer it shortly and directly without repeating the question.‘
We include audio information by adding the transcript in the prompt, in the following format: ‘[START TIMESTAMP] - [END TIMESTAMP]: [SUBTITLE]‘. In case of modality ablation, we simply remove either the subtitles or the video frames from the inputs and adapt the query accordingly.
| Multiple-Choice QA | ||||||
| V | L | VL | ||||
| FrozenBiLM | ||||||
| Shot-level | 22.1 | 30.9 | 30.8 | |||
| Scene-level | 22.7 | +0.6 | 34.8 | +3.9 | 34.8 | +4.0 |
| Movie-level | 23.4 | +1.3 | 38.2 | +7.3 | 38.6 | +7.8 |
| Video-Llava-7B | ||||||
| Shot-level | 16.0 | 24.4 | 24.0 | |||
| Scene-level | 23.7 | +7.7 | 20.1 | -4.3 | 20.1 | -3.9 |
| Movie-level | 34.2 | +18.2 | 21.3 | -3.1 | 24.7 | +0.7 |
| Llovi | ||||||
| Shot-level | 32.8 | 38.5 | 50.1 | |||
| Scene-level | 34.2 | +1.4 | 51.2 | +12.7 | 55.4 | +5.3 |
| Movie-level | 30.8 | -2.0 | 64.2 | +25.7 | 55.6 | +5.5 |
| Human | ||||||
| Movie-level | 59.0 | 70.9 | 89.8 | |||
Metrics. For multiple-choice QA, we use regular expression to extract the digit choice inside the parenthesis in the model’s response. We then compute the overall accuracy. For open-ended QA, we use a LLM to compare the free-form response from the model to the correct answer. In details, we prompt GPT-3.5 (‘gpt-3.5-turbo-0125‘) for the OpenAI API, and use the prompt provided by [39]. We get a yes/no value and an associated score. We then compute the average score and the overall accuracy.
Appendix F Failure cases analysis
In this section, we provide an analysis of failure cases to better understand the limitations of our model in video question answering. Table 12 categorizes these failure cases into three key domains: visual cues, where the model struggles to interpret visual details critical to answer the question; character identification, where the model fails to correctly associate characters with their names, actions, or attributes; and temporal understanding, which highlights challenges in reasoning about events over time. By examining these errors, we aim to emphasize specific weaknesses and inform future improvements in multimodal reasoning and long-term video understanding.
Appendix G Technical details on the temporal window study
In this section, we present additional details and results for the temporal window study (see Section 4.5). The goal of this experiment is to evaluate the performance of various methods on SF20K-Test at different temporal levels—shot, scene, and movie—across all modalities. We examine FrozenBiLM [6], LLoVi [83], and Video-Llava-7B [36].
At the shot-level, we truncate the input prompt to include only information from a specific shot. For FrozenBiLM and Video-LLaVA, it means that we only include subtitles and frames extracted from the shot. For LLoVi and LangRepo, we only include subtitles and captions extracted from the shot. Hence, we get logits for each answer option and shot, aggregated in a matrix of shape : we obtain a final prediction by taking the maximum logit over both dimensions.
The scene-level inference is similar to the shot-level one, the main difference being that we constrain the input information (subtitles and frames/captions) to the scene-level, which arbitrarily corresponds to 10 shots in our case. We end up with a logit matrix of size . The movie-level inference is straightforward: it is the same as in Section 4.3.
The numerical results of the temporal window study are reported in Table 13. Along the temporal direction (i.e. row-wise for each method), we observe that: (i) generally, all methods exhibit consistent gains with larger temporal windows confirming the importance of long-term understanding for our designed questions, and (ii) exceptions are noted, such as with Video-Llava-7B in Language and Vision-Language settings, likely due to its low precision (near random chance at 20%) which diminishes the impact of the temporal window.
Appendix H Prompts